# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 2.5. Application to compact pp-waves
The so-called pp-waves are a special class of Lorentz- ian manifolds with a parallel null vector field $V$, with a great importance in General Relativity. It was proved in [17] that a compact pp-wave is complete. The completeness of the leaves of the codimension $1$ totally geodesic foliation $\mathfrak{F}$ tangent to $V^{\perp}$ was a crucial step in the proof. On the other hand, it is observed in Example 2.4 that the leaves of $\mathfrak{F}$ have a natural $(\mathsf{L}_{\sf{u}}(k),\mathbb{R}^{k+1})$-structure. We then recover, applying Corollary 2.16, the completeness result along the leaves of the codimension $1$ foliation $\mathfrak{F}$. We see that the completeness of the $\mathfrak{F}$-leaves of a pp-wave is independent of the presence of a Lorentzian ambient metric, and is therefore purely a leafwise completeness problem!
@@ 3. $(G,X)$-foliated completeness problem: affine lightlike geometry
Let $(M,l)$ be a “weakly pp-wave” of dimension $n+2$, that is, a Lorentzian manifold with a null parallel line field $l$, for which the (totally geodesic) foliation $\mathfrak{F}$ orthogonal to $l$ is flat with respect to the induced connection. Then, the leaves of $\mathfrak{F}$ have a natural affine lightlike geometry. So weakly pp-waves give natural examples of $(\mathsf{L}(n),\mathbb{R}^{n+1})$-foliated manifolds.
We ask whether a completeness statement analogous to Theorem 2.5 still holds for this geometric structure $({\sf{L}}(n),\mathbb{R}^{n+1})$. Notice that the methods from the previous case of $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry give straightforwardly that the developing map $D$ (on the universal cover of some leaf) is injective, and surjective on $\mathcal{F}/\mathcal{V}_{M}$, but fail in the surjectivity along the $\mathcal{V}_{M}$-leaves. This failure is due to lack of an invariant vector field. Indeed, the key point for surjectivity along the $\mathcal{V}_{M}$-leaves was that the $V_{M}$-flow is equivariant and complete. Formulating this observation we get the following
@@
However, it turns out that the incomplete situation can occur. In the following we give some examples illustrating this. We present two kind of examples: the first kind is an affine manifold with a codimension $1$ foliation that supports an $({\sf{L}}(n),\mathbb{R}^{n+1})$-structure, and the second kind is a “weakly pp-wave”.
@@
Here, $\Gamma$ is a subgroup of $\sf{Aff}(\mathbb{R}^{3})$ acting properly discontinuously and freely on $U$, with a compact quotient. So $\Gamma\backslash U$ is a compact affine manifold: the product of a Euclidean torus and an Hopf circle of affine length $\lambda$. Now, equip $\mathbb{R}^{3}$ with the Lorentzian flat metric $2dudv+dx^{2}$. The subgroup $\Gamma$ does not preserve the flat metric, but it preserves the null direction $\mathbb{R}V$, where $V=\partial_{v}$, together with the degenerate Riemannian metric $dx^{2}$ on the leaves of $V^{\perp}$. Hence the foliation tangent to $V^{\perp}$ is a $({\sf{L}}(1),\mathbb{R}^{2})$-foliation on $\Gamma\backslash U$. Note that if $\Gamma$ is a subgroup of the Poincaré group, i.e. preserves the Lorentzian flat metric, then $U$ is necessarily $\mathbb{R}^{3}$ and the quotient is complete due to Carriére’s result. We see that in this case, having a Lorentzian ambient structure is not artificial regarding the foliated-completeness question.
@@
Let $(T_{CP},g):=(\mathbb{R}^{2}\smallsetminus\{(0,0)\},\frac{2dxdy}{x^{2}+y^{2}})/(x,y)\sim 2(x,y)$ denote the Clifton-Pohl torus. Let $\mathcal{N}$ be one of its two null foliations. Then the tangent bundle of $\mathcal{N}$ defines a parallel null line on $T_{CP}$. It is known that the leaves of the foliation are incomplete.
Example 3.5 (Higher dimension).
@@
(1) Consider the product $M=T_{CP}\times\mathbb{S}^{1}$, with Lorentzian metric $g+dt^{2}$. Denote by $\mathcal{N}$ the null foliation on $M$ coming from one of the null foliations of $T_{CP}$. Its tangent bundle defines a null parallel line field on $M$. The orthogonal leaves $\mathcal{F}=\mathcal{N}^{\perp}$ are given by $\mathcal{N}\times\mathbb{S}^{1}$. They are flat (with respect to the induced connection) and incomplete. In this example, a leaf of $\mathfrak{F}$ is either a cylinder or a torus, and the universal cover of a leaf develops onto ×$\mathbb{R}\times I$, where $I$ is some open interval in $\mathbb{R}$ which is either affinely equivalent to ]0,1[]0,1[] 0 , 1 [ or to ]0,+[]0,$\infty$[] 0 , + ∞ [ (both cases occur).
(2) One can deform the previous example in order to have simply connected $\mathfrak{F}$-leaves. Consider $X=\mathbb{R}^{3}\smallsetminus\{(0,0\}\times\mathbb{R}$ with Lorentzian metric $g+dt^{2}$. Its isometry group is $G=\mathbb{R}\times\mathbb{R}$ (up to finite index), where the first $\mathbb{R}$-factor acts by a $1$-parameter group of homotheties on the Clifton-Pohl metric of $\mathbb{R}^{2}\smallsetminus\{(0,0)\}$, and the second $\mathbb{R}$-factor acts by translation on $\mathbb{R}$. The group $G$ preserves the foliation $\mathcal{N}\times\mathbb{R}$ of $X$. Consider a discrete subgroup of $G$ acting properly cocompactly on $X$, and not intersecting the second $\mathbb{R}$-factor. The quotient is a Lorentzian torus with a parallel null line field, and having simply connected orthogonal leaves.
[…]
@@
In the above examples, incompleteness occurs along geodesics tangent to $\mathcal{N}$, but also on spacelike geodesics not tangent to $\mathbb{S}^{1}$ (a flat band of the form ×$\mathbb{R}\times]0,1[\subset\mathbb{R}^{2}$ contains spacelike lines that leave the band). However, in this example, any spacelike geodesic tangent to a leaf is contained in an infinite band (infinite here is for the Riemannian transverse structure).
4. Foliated $({\sf{L}}_{\sf{u}}(k,1),\mathbb{R}^{k+2})$-geometry and its dual
In this section we show that the foliated completeness Problem 1.8 is not true for the $({\sf{L}}_{\sf{u}}(1,k),\mathbb{R}^{k+2})$-structure and its dual, by constructing concrete incomplete examples.
4.1. Compact $({\sf{L}}_{\sf{u}}(k,1),\mathbb{R}^{k+2})$-manifolds
In a recent paper [3], it is shown that a compact $({\sf{L}}_{\sf{u}}(k,1),\mathbb{R}^{k+2})$-manifold is complete. The proof uses techniques from [7, 9, 6].
4.2. Compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifolds
@@
The dual structures were mentioned earlier by Carrière in [6, Theorem 3.2.1]. Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifold. This structure on $M$ gives rise naturally to a codimension $1$ totally geodesic Lorentzian foliation with flat leaves. In [6, Theorem 3.2.1], it was shown that the developing map on those leaves is injective. Here we use the extra invariant $1$-form to show that the developing map of $M$ is itself injective.
@@
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifold. The developing map $d$ is injective. As a result, the holonomy representation is injective, and the fundamental group is isomorphic to a discrete subgroup of ${\sf{L}}_{\sf{u}}^{*}(k,1)$.
@@
The proof for the injectivity of $d$ is the same as in Corollary 2.17. The rest is an immediate consequence. ∎
This allows to show
@@
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(k,1),\mathbb{R}^{k+2})$-manifold. If $\pi_{1}(M)$ is virtually solvable, then $M$ is complete.
@@
First, we remark that $M$ is aspherical. Indeed, $\widetilde{M}$ fibers over $\mathbb{R}$ with a contractible fiber (corresponding to the Lorentzian flat leaves, see [6, Proposition 3.2.2]). Moreover, since $\pi_{1}(M)$ is virtually isomorphic to a discrete subgroup of a connected Lie group, $\pi_{1}(M)$ is virtually solvable if and only if it is virtually polycyclic [12, Section 1]. Hence, if $\pi_{1}(M)$ is virtually solvable we have $\mathsf{Rank}(\pi_{1})=\mathsf{cd}(\pi_{1})$=$\dim(M)$. However, $M$ has a parallel volume. The claim follows using [11, Theorem 3.5]. ∎
@@
Fact 4.1 implies that the holonomy is injective. Hence, the $\pi_{1}(M)$ is isomorphic to a subgroup of ${\sf{L}}_{\sf{u}}^{*}(1,1)$ Clearly, ${\sf{L}}_{\sf{u}}^{*}(1,1)$ is solvable. So, the claim follows from Corollary 4.2. ∎

### v2

@@ 2.5. Applications
The special class of Lorentzian manifolds, known as pp-waves, play a significant role in General Relativity. It was proved in [23] that a compact pp-wave is complete. A key step in the proof was establishing the completeness of the leaves of the codimension $1$ totally geodesic foliation $\mathfrak{F}$ tangent to $V^{\perp}$. As observed in Example 1.9, the leaves of $\mathfrak{F}$ inherit a natural $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-structure, compatible with the connection induced from the Lorentzian metric. We then recover, applying Corollary 2.20, the completeness result along the leaves of $\mathfrak{F}$. We see that the completeness of the $\mathfrak{F}$-leaves of a pp-wave is independent of the presence of a Lorentzian ambient metric, and is, in fact, a purely leafwise completeness phenomenon!
2.5.2. Compact manifolds modeled on (non-flat) homogeneous plane waves
Within the family of pp-waves, there exists a distinguished subclass of metrics known as plane waves, defined by specific conditions on the Riemann curvature tensor. These metrics are widely studied due to their very interesting mathematical properties.
By a compact model of a homogeneous plane wave, we mean a compact $(G_{\rho},X_{\rho})$-manifold, where $X_{\rho}$ is a non-flat simply connected homogeneous plane wave, and $G_{\rho}$ is the identity component of its isometry group. These groups are described in [17]. For $\dim X_{\rho}=n+2$, we have
$G_{\rho}=(\mathbb{R}\times K)\ltimes{\mathsf{Heis}},\;\;X_{\rho}=G_{\rho}/I,\,I\subset K\ltimes{\mathsf{Heis}},$
where $K\ltimes{\mathsf{Heis}}\subset\mathsf{L}_{\sf{u}}(n)$. We denote by $V$ the parallel null vector field of $X_{\rho}$, and by $\mathfrak{F}$ the foliation tangent to $V^{\perp}$. The subgroup of $G_{\rho}$ fixing a leaf of $\mathfrak{F}$ is exactly $K\ltimes{\mathsf{Heis}}$.
Non-flat homogeneous plane waves are divided into two families: one where the model space $X_{\rho}$ is geodesically complete and the parallel null vector field $V$ is preserved by $G_{\rho}$, and another where $X_{\rho}$ is incomplete and only the line field $\mathbb{R}V$ is preserved by $G_{\rho}$.
Theorem 2.23.
A compact manifold $M$ modeled on a non-flat simply connected homogeneous plane wave $(G_{\rho},X_{\rho})$ is $(G_{\rho},X_{\rho})$-complete.
Proof.
If $M$ is modeled on the first family, it admits a plane wave structure. To see this, consider the pullback of $V$ and the Lorentzian metric $g$ on the model space $X_{\rho}$ by the developing map $\mathsf{Dev}:\widetilde{M}\to X_{\rho}$. Since both $V$ and $g$ are invariant under the action of $G_{\rho}$, this induces a Lorentzian metric and a vector field on the universal cover $\widetilde{M}$, both invariant under the deck transformations by $\pi_{1}(M)$. Thus, $M$ inherits a Lorentzian structure, with a parallel null vector field, locally isometric to that of $X_{\rho}$, establishing that $M$ is a plane wave. This manifold is complete by [23, 26]. In contrast, if $M$ is modeled on the second family, it is not a plane wave but rather a weakly plane wave. It has a parallel null line field $l$ instead of a parallel null vector field. The distribution $l^{\perp}$ is tangent to a lightlike totally geodesic foliation $\mathfrak{F}_{M}$. Even if $V$ is not preserved by the whole group $G_{\rho}$, it appears from the description of $G_{\rho}$ that the subgroup fixing a leaf is a subgroup of the affine unimodular group ${\sf{L}}_{\sf{u}}(n)$,  $n+2=\dim X_{\rho}$. Thus, the holonomy representation of an $\mathfrak{F}$-leaf takes values in ${\sf{L}}_{\sf{u}}(n)$. Consequently, $(M,\mathfrak{F}_{M})$ is an $({\sf{L}}_{\sf{u}}(n),\mathbb{R}^{n+1})$-foliated manifold. We claim that $M$ is $(G_{\rho},X_{\rho})$-complete, despite being geodesically incomplete (the geodesic incompleteness comes from the model space!). Indeed, by Theorem 2.10, each leaf in the universal cover $\widetilde{M}$ is mapped bijectively to a leaf in the model via the developing map. On the other hand, there is a $G_{\rho}$-invariant closed $1$-form on $X_{\rho}$, whose kernel is equal to $V^{\perp}$. Indeed, the space of $\mathfrak{F}$-leaves is affinely identified with the $\mathbb{R}$-factor in $G_{\rho}$, on which $G_{\rho}$ acts by affine transformations: $K\ltimes{\mathsf{Heis}}$ acts trivially, while the $\mathbb{R}$-factor acts by translation. As a result, there is a closed $G_{\rho}$-invariant $1$-from $\omega_{0}$ on the space of leaves. Pulling back $\omega_{0}$ via the projection map $X_{\rho}\to X_{\rho}/\mathfrak{F}$ yields the desired $1$-form on $X_{\rho}$. As in the proof of Corollary 2.22, this allows to define a flow on $\widetilde{M}$ that acts transversely on the foliation, mapping leaves to one another equivariantly with respect to the developing map. We conclude as in the proof of Corollary 2.22. ∎
@@ 3. $(G,X)$-foliated completeness problem: affine lightlike geometry
We now ask whether a completeness result analogous to Theorem 2.10 holds for this geometric structure $({\sf{L}}(n),\mathbb{R}^{n+1})$. The methods from the previous case of the $(\mathsf{L}_{\sf{u}}(n),\mathbb{R}^{n+1})$-geometry readily show that the developing map ${\mathsf{Dev}}:\widetilde{F}\to\mathbb{R}^{n+1}$ on the universal cover of some leaf $F$ is injective, and surjective from $\widetilde{F}/\widetilde{\mathcal{V}}_{M}$ onto the space of $\mathcal{V}$-leaves. However, surjectivity may fail along the $\mathcal{V}_{M}$-leaves. This failure is due to the absence of an invariant vector field. Indeed, the key ingredient for surjectivity along the $\mathcal{V}_{M}$-leaves was the existence of a vector field $V_{M}$, whose flow was both equivariant and complete. Formulating this observation we get the following
@@
It turns out that the incomplete situation can occur. In what follows, we give some examples illustrating this. We present two kind of examples: the first kind is an affine manifold with a codimension $1$ foliation that supports an $({\sf{L}}(n),\mathbb{R}^{n+1})$-structure, and the second kind is a “weakly pp-wave”.
@@
Here, $\Gamma$ is a subgroup of $\sf{Aff}(\mathbb{R}^{3})$ acting properly discontinuously and freely on $U$, with a compact quotient. So $\Gamma\backslash U$ is a compact affine manifold. It is the product of a Euclidean torus and an Hopf circle of affine length $\lambda$. Now, equip $\mathbb{R}^{3}$ with the Lorentzian flat metric $2dudv+dx^{2}$. The group $\Gamma$ does not preserve the metric, but it preserves the null direction $\mathbb{R}V$, where $V=\partial_{v}$, together with the degenerate Riemannian metric $dx^{2}$ on the leaves of $V^{\perp}$. Hence, the foliation tangent to $V^{\perp}$ induces a $({\sf{L}}(1),\mathbb{R}^{2})$-foliation on $\Gamma\backslash U$. Note that if $\Gamma$ is a subgroup of the Poincaré group, i.e. preserves the Lorentzian flat metric, then $U$ is necessarily $\mathbb{R}^{3}$ and the quotient is complete due to Carrière’s result. We see that in this case, having a Lorentzian ambient structure is not artificial regarding the foliated completeness question.
@@
Let $(T_{CP},\bar{g}):=(\mathbb{R}^{2}\smallsetminus\{(0,0)\},g:=\frac{2dxdy}{x^{2}+y^{2}})/(x,y)\sim 2(x,y)$ denote the Clifton-Pohl torus. Let $\mathcal{N}$ be one of its two null foliations. Then, the tangent bundle of $\mathcal{N}$ defines a parallel null line field on $T_{CP}$. It is known that the leaves of the foliation are incomplete. And $\mathcal{N}$ has a natural affine structure, which is in fact a homothety structure $(\mathbb{R}^{*}_{+},\mathbb{R})$.
Example 3.4 (Higher dimensions).
@@
(1) Consider the product $M=T_{CP}\times\mathbb{S}^{1}$, with Lorentzian metric $\bar{g}+\overline{dt^{2}}$. Denote again by $\mathcal{N}$ one of the null foliations of $T_{CP}$. Consider the lightlike and totally geodesic foliation $\mathfrak{F}:=\mathcal{N}\times\mathbb{S}^{1}$ of $M$. The line field $T\mathfrak{F}^{\perp}$ on $M$ is null and parallel. The leaves of $\mathfrak{F}$ are flat (with respect to the induced connection) and incomplete. So $M$ is a weakly pp-wave having incomplete leaves. A leaf of $\mathfrak{F}$ is either a cylinder or a torus. And the universal cover of a leaf develops onto ×$\mathbb{R}\times I$, where $I$ is some open interval in $\mathbb{R}$ which is either affinely equivalent to ]0,1[]0,1[] 0 , 1 [ or to ]0,+[]0,$\infty$[] 0 , + ∞ [ (both cases occur).
@@
Incompleteness occurs exactly along the null geodesics tangent to $\mathcal{N}\times\mathbb{S}^{1}$, and along spacelike geodesics not tangent to the $\mathbb{S}^{1}$-factor (a flat band of the form ×$\mathbb{R}\times I\subset\mathbb{R}^{2}$, with $I$ as above, contains spacelike lines that leave the band).
(2) The previous example can be modified in order to have all $\mathfrak{F}$-leaves non-compact. Consider $X=\mathbb{R}^{2}\smallsetminus\{(0,0)\}\times\mathbb{R}$ with Lorentzian metric $g+dt^{2}$. Its isometry group, up to finite index, is $G=\mathbb{R}\times\mathbb{R}$, where the first $\mathbb{R}$-factor acts by a $1$-parameter group of homotheties on $(\mathbb{R}^{2}\smallsetminus\{(0,0)\},g)$, and the second $\mathbb{R}$-factor acts by translation on $(\mathbb{R},dt^{2})$. The group $G$ preserves the foliation $\mathcal{N}\times\mathbb{R}$ of $X$. Now, consider a discrete subgroup $\Gamma\subset G$ acting properly freely and cocompactly on $X$, and not intersecting the second $\mathbb{R}$-factor. The quotient $\Gamma\backslash X$ is a Lorentzian torus, where the induced foliation consists of leaves that are either planes or cylinders. By Remark 2.9, one cannot find a quotient space where all the leaves of the foliation are simply connected.
(3) Taking the product of $T_{CP}$ with a compact Riemannian manifold (of higher dimension) gives an example with non-flat $\mathfrak{F}$-leaves.
4. Pseudo-Riemannian variants and their duals
In this section, we will see that leafwise completeness may fail for the $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-structure and its dual. We will construct explicit examples of incomplete foliations. But before that, we first consider the case of a codimension 00 foliation, i.e. when the structure is defined on a compact manifold. We restrict to the case $q=1$.
4.1. Compact $({\sf{L}}_{\sf{u}}(n,1),\mathbb{R}^{n+2})$-manifolds
Completeness of compact $({\sf{L}}_{\sf{u}}(n,1),\mathbb{R}^{n+2})$-manifolds is an open question and a natural part of the Markus conjecture, as mentioned in the introduction. We also note that the method in [6] does not apply here, even for $n=1$, due to the fact that the discompacity of ${\sf{L}}_{\sf{u}}(n,1)$ is $2$.
4.2. Compact $({\sf{L}}_{\sf{u}}^{*}(n,1),\mathbb{R}^{n+2})$-manifolds
@@
The dual structures were mentioned earlier by Carrière in [6, Theorem 3.2.1]. Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(n,1),\mathbb{R}^{n+2})$-manifold. This structure on $M$ gives rise naturally to a codimension $1$ totally geodesic Lorentzian foliation $\mathfrak{S}$ with flat leaves. In [6, Theorem 3.2.1], it is shown that the developing map is injective on each $\mathfrak{S}$-leaf. Here, we can use the extra invariant $1$-form to show that the developing map of $M$ is itself injective.
@@
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(n,1),\mathbb{R}^{n+2})$-manifold. The developing map ${\mathsf{Dev}}:\widetilde{M}\to\mathbb{R}^{n+2}$ is injective. As a result, the holonomy representation is injective, and the fundamental group is isomorphic to a discrete subgroup of ${\sf{L}}_{\sf{u}}^{*}(n,1)$ acting properly freely on ${\mathsf{Dev}}(\widetilde{M})$.
@@
We already know that ${\mathsf{Dev}}$ is injective on a $\mathfrak{S}$-leaf ([6, Theorem 3.2.1]). Global injectivity of ${\mathsf{Dev}}$ follows identically to the second part of the proof of Corollary 2.22. The rest is an immediate consequence. ∎
@@
Let $M$ be a compact $({\sf{L}}_{\sf{u}}^{*}(n,1),\mathbb{R}^{n+2})$-manifold with a virtually solvable holonomy group. Then $M$ is complete.
@@
First, observe that $M$ is aspherical. Indeed, $\widetilde{M}$ fibers over $\mathbb{R}$ with a contractible fiber (the fibers are the leaves of $\mathfrak{S}$, which are contractible by [6, Proposition 3.2.2]). This yields $\mathsf{cd}(\pi_{1}(M))=\dim(M)$, where $\mathsf{cd}(\pi_{1}(M))$ is the cohomological dimension of $\pi_{1}(M)$. On the other hand, by Fact 4.1 we get that $\pi_{1}(M)$ injects into a discrete subgroup of the connected Lie group ${\sf{L}}_{\sf{u}}^{*}(n,1)$. However, by assumption $\pi_{1}(M)$ is virtually solvable, then by [28, Lemma 2.1] $\pi_{1}(M)$ is even virtually polycyclic. In particular, we get $\mathsf{Rank}(\pi_{1}(M))=\mathsf{cd}(\pi_{1}(M))=\dim(M)$. Since $M$ admits a parallel volume form, the claim follows from [15, Corollary 3.8]. ∎
@@
Fact 4.1 implies that the holonomy representation is injective. Hence, $\pi_{1}(M)$ is isomorphic to a subgroup of ${\sf{L}}_{\sf{u}}^{*}(1,1)$ Clearly, ${\sf{L}}_{\sf{u}}^{*}(1,1)$ is solvable. So, the claim follows from Corollary 4.2. ∎